# Robust $\mathcal{H}_{\infty}$ control and worst-case search in constrained parametric space

### v4

@@ V-B Closed loop and control problem
The benchmark focuses on the attitude control of the satellite described above to follow a reference $r$. The control input $u$ of the plant $G(s,\delta)$ is the torque applied to the satellite (e.g. by reaction wheels, thrusters…) and the output vector contains the 3 angle measurements polluted by some noise $n$. The closed loop is represented in Figure 2, where $d(s)$ represents avionics: a 2-nd order Pade approximation of a 0.125s delay, and a 2-nd order low-pass filter with a cut-off frequency of 10 rad/s representing actuator dynamics. We note $\mathcal{K}$ the set of controllers $K(s)$ of the desired structure: one 4-th order controller per spacecraft axis, each one taking the measured angle and rate as inputs and generating a control torque, resulting in 24 decision parameters per controller (72 in total). The closed-loop plant has 50 states.
@@
The control problem consists in placing a control bandwidth of 0.1 rad/s (hence interacting with the uncertain flexible modes) while ensuring a modulus margin of 0.5 and minimizing the variance of the actuator command in response to measurement noise. It reads:
@@ V-C Worst-case search
In this section, we present the worst-case search for the spectral abscissa and the $\mathcal{H}_{\infty}$ norm, as presented in Section III. We selected three controllers $K_{1}$, $K_{2}$, $K_{3}$ obtained at early iterations of the robust controller tuning presented in Section V-D and analyze the robustness of the closed-loop system with three Test cases. For the sake of brevity, and since all 3 test cases yield similar conclusions, we present only the Test Case 1 in the body of the paper; Test Cases 2 and 3 are in Appendix B. Test case 1 illustrates how the proposed approaches perform in presence of several local optima, in different regions of the parametric space, that are not particularly rare. We search for $\max_{\delta\in\mathcal{D}}\;\alpha(A(K_{1},\delta))$.
@@
We ran the Sequential Quadratic Programming (SQP) algorithm from 100 random initial points. As termination criteria, we set $10^{-6}$ as optimality tolerance (KKT stationarity condition), and $10^{-8}$ as step tolerance (step size between two iterations). The resulting worst-case spectral abscissas $\alpha(A(\delta^{*}_{\text{SQP}}))=-a(\delta^{*}_{\text{SQP}})$ of each run are in Figure 3(a), and suggest the presence of several local optima, which highlights the limitation of such a local method. These various local optima may correspond to significantly different parametric configurations. For example, the worst case 0.135 has the value -1.0 for the (normalized) value of one of the moments of inertia of the solar panel, -0.09 and 1.0 for the modal participation factors of the clamped-free modes 1 and 3 in rotation around the same axis; in contrast, the local optimum 0.007 has -0.30, -1.0 and -0.87 respectively for these three parameters. Figure 3(b) presents the poles map for the initial poles (for the random initialization of the parameters) and final poles (corresponding to the worst case) for one of the runs that converged to the suggested global optimum.
Excluding the runs that did not detect instability (each of which took less than 0.1s), the remaining runs took between 2.5 and 78s (average: 21s – regarding total computation time, we recall that each SQP run is inherently sequential, but that the 100 runs are executed in parallel).
@@
(b) Poles map (zoomed) for a SQP run
(c) PSO and PSO+SQP, with random 100 initializations
(d) MC and MC+SQP, repeated 100 times
@@
b) PSO, PSO + SQP
The parameters of the Particle Swarm Optimization (PSO) were chosen as follows to favor global exploration over local exploitation. We set a range of 0.6 to 1.1 for the adaptive inertias (’InertiaRange’ option in MATLAB). Iterations terminate when performance is increased by less than 10% (”FunctionTolerance’) across 20 last iterations (’MaxStallIterations’). For a given computational budget, because PSO is sensitive to initialization [35], running several independent instances with different initializations can yield better results than increasing the number of particles or tightening the termination tolerance. Therefore, we performed 100 independent runs with the parameters given above, each with 500 particles (’SwarmSize’) – cf. also [43] for a discussion on population size in PSO. All other parameters were set to MATLAB’s default settings. Bearing in mind that our parameter choices were specifically tuned for multiple fast independent runs rather than to maximize the reliability of a single run, the results should be interpreted to determine how many such runs are required to reliably identify a worst case. We see that, with a few runs, we are able to consistently identify the worst-case configuration that was most often missed by the local SQP alone.
For each run, the computations of the 500 particles were parallelized over 14 cores. Each run took between 6s and 8s (average: 7s) with 11000 model evaluations. Each run was followed by SQP (starting from the identified worst case), each between 37 and 97s (average: 77s) ; these SQPs were parallelized. The results in Figure 3(c) show that the (suggested) global optimum can be consistently found using a few random initializations of PSO+SQP.
c) MC, MC + SQP
To enable a fair comparison with PSO, (i) we performed a total of $N_{\text{MC}}=$1.1e6 evaluations of the objective function $a$, matching the total number of objective function evaluations of the PSO across all 100 runs, and (ii) we divided the $N_{\text{MC}}$ random samples into 100 groups, that we call Monte-Carlo (MC) experiments. MC performed almost as well as PSO, cf. Figure 3(d). In all runs, the subsequent SQP significantly improves the result. Each MC run took about 6.2s, and the SQP runtimes were similar to the PSO case.
We mention that 78% of the drawn samples were not admissible ($c(\delta)>0$) and were therefore discarded before evaluation of $a(\delta)$. About 92% of the admissible configurations were destabilizing ($a(\delta)<0)$. We also mention that the total number of runs $N_{\text{MC}}$ verifies $N_{\text{MC}}\approx\textrm{ln}(\gamma)/\textrm{ln}(1-\epsilon)$ with the values $\gamma=\epsilon=$1e-5, which guarantees as in equation (21) that the worst-case $a(\delta^{*}_{\text{MC}})$ found through all samples of Figure 3(d) (without considering the subsequent SQP) covers 99.999% of the probability distribution of $a(\delta)$ with a confidence level of 99.999%. These observations suggest that the worst case around 0.136 is rare, even though unstable configurations are frequent in Test case 1.
@@ V-D1 Implementation details
Compared to the choice of parameters of Section V-C, here we reduced the swarm size to 100 and we relaxed the PSO convergence criterion: iterations terminate when performance is increased by less than 10% across 10 last iterations. For SQP, we reduced the step tolerance from 1e-8 to 1e-6. The objective is to speed up the detection of active configurations in most iterations, relying on Step 4 to perform a finer validation. Below, we present three different strategies for the worst-case searches.
Strategy 1
We run a single instance of PSO, followed by SQP. Although, as seen in Section V-C, this does not always converge to the global optimum, we prioritize computational speed by stopping at the first KKT point for several reasons. First, each KKT point may be valuable to add to the active configurations even if it is not the global worst-case for the current controller. Second, because multiple iterations are performed, any missed problematic configuration can be detected in subsequent iterations. Third, if stability is validated in Step 2 but a destabilizing configuration is discovered when degrading performance in Step 3, the process returns to Step 2. Fourth, a more exhaustive search is carried out in Step 4, allowing a return to previous iterations if a problematic configuration is identified, while limiting the number of such more expensive searches. The resulting (single) KKT point, if problematic, is added to the active configurations.
Strategy 2
Strategy 2 is the same as Strategy 1, except that the PSO is stopped as soon as a problematic configuration is found (using ’ObjectiveLimit’ option in MATLAB); the PSO is then followed by SQP.
Strategy 3
In Strategy 3, we perform 4 instances of PSO+SQP and obtain 4 KKT points $\delta^{*}_{i}$. For clarity, let us assume that they are all problematic configurations. The worst case among those 4, noted $\delta^{*}_{1}$, is added to the active configurations. Then, we compute the distance $|\delta_{2}^{*}-\delta_{1}^{*}|$; if the distance exceeds a specified threshold, $\delta^{*}_{2}$ is also added to the active configurations. The procedure is repeated for $\delta_{3}^{*}$ and $\delta_{4}^{*}$: each $\delta_{i}^{*}$ is added to the active scenarios only if its distance from all previously selected points exceeds the threshold. The threshold is chosen as half of the hypercube diagonal of the parameter space, i.e., $\sqrt{k}$, where $k$ is the number of uncertain parameters. This choice limits the number of active scenarios by enforcing a minimum spatial separation, thereby ensuring that only significantly distinct worst-case configurations are retained. Indeed, the computation time of the multi-models controller synthesis (Step 1) increases with the number of active configurations.
@@
In Step 4, we run up to $N$ instances of PSO. If a problematic configuration is found, we run SQP, and add the resulting worst case to the active configurations. Otherwise, we run SQP on all $N$ instances. In other words, Step 4 operates like Section V-C when no problematic configuration is found, but terminates earlier when one is identified. This procedure is applied for both the stability and performance, and the Algorithm 2 terminates when no problematic configuration is found. In the implementation, we conducted the worst-case search for performance before that for stability to avoid unnecessary computations, since performance was observed to fail more frequently than stability. We used the same optimization parameters as in Section V-C and $N=20$, which is deemed sufficient to ensure thorough validation of the robustness.
@@ V-D2 Results
As in section V-C, the tests are ran on a 12th Gen Intel(R) Core(TM) i7-12700H 2.30 GHz, with 14 cores and 16GB of RAM, and computation times are provided for reference. We used the tolerances $\epsilon_{1}=5\%$ and $\epsilon_{2}=1\%$ (cf. equation (23)).
The Algorithm 2 was executed 5 times for each strategy. All runs converged to a robust controller. Table I presents a few averaged metrics for comparison. The runtime of the final validation (Step 4) is averaged over all 15 runs, because this step is conducted identically for each strategy. As the multi-models controller synthesis (Step 1) becomes increasingly long when the number of active configurations increases, Strategy 2 results in the longest runtime. Indeed, although it saves time by prematurely stopping the PSO, the resulting KKT point is typically not a global optimum, which in turn requires a larger number of active configurations to ensure robustness. Strategy 3, which may add multiple KKT points to the active configurations at each iteration, results in less iterations and a reduced total runtime.
Figure 4 presents the transfer functions of Problem (26) after optimization of the controller. Active configurations (in red) include the worst cases detected in the final iteration.
@@
Metrics Strategy 1 Strategy 2 Strategy 3
Nb of iterations ( = nb. of executions of Step 1) (min/avg/max) 7/8.4/10 11/12.2/17 6/7.4/10
@@
Total runtime, excluding final validation (avg) 4034s 4919s 3708s
… including % represented by controller tuning (Step 1) 34% 58% 33%
… including % represented by worst-case searches (Steps 2 and 3) 66% 42% 67%
@@ VI Conclusion
This paper focused on the robust control and worst-case search in presence of nonlinear constraints restricting the hypercube. This specificity is missing from the current state of the art, and arises for example when considering uncertain mechanical properties e.g. in flexible satellites.
@@
To perform the worst-case search, we combined Particle Swarm Optimization (or Monte-Carlo sampling, if stochastic guarantees are desired) for global exploration and Sequential Quadratic Programming for local exploitation. Numerical experiments on a flexible spacecraft benchmark show that this method is able to reliably detect even rare worst-case configurations. Then, by iteratively constructing a set of active configurations, multi-objectives robust controller optimization was performed in reasonable runtime.
Finally, we emphasize that the proposed worst-case search, performing global exploration (albeit without deterministic guarantee of finding the global optimum) and being compatible with a large number of uncertain parameters and states (as a gradient-based method), has the potential to complement existing robust control techniques even in the unconstrained case.

### v5

@@ V-B Closed loop and control problem
The benchmark focuses on the attitude control of the satellite described above to follow a reference $r$. The control input $u$ of the plant $G(s,\delta)$ is the torque applied to the satellite, and the output vector contains the 3 angle measurements polluted by some noise $n$. The closed loop is represented in Fig. 2, where $d(s)$ represents actuator dynamics. We note $\mathcal{K}$ the set of controllers $K(s)$ of the desired structure: one 4-th order controller per spacecraft axis, each one taking the measured angle and rate as inputs and generating a control torque, resulting in 24 decision parameters per controller (72 in total). The closed-loop plant has 50 states.
@@
The control problem consists in placing a control bandwidth of 0.1 rad/s (hence interacting with the uncertain flexible modes and causing a strong sensitivity to the uncertain parameters) while ensuring a modulus margin of 0.5 and minimizing the variance of the actuator command in response to measurement noise. It reads:
@@ V-C Worst-case search
In this section, we present the worst-case search for the spectral abscissa and the $\mathcal{H}_{\infty}$ norm, as presented in Section III. We selected three controllers $K_{1}$, $K_{2}$, $K_{3}$ obtained at early iterations of the robust controller tuning presented in Section V-D and analyze the robustness of the closed-loop system with three Test cases. For the sake of brevity, and since all 3 Test cases yield similar conclusions, we present only the Test case 1 in the body of the paper; Test cases 2 and 3 are in Appendix B. Test case 1 illustrates how the proposed approaches perform in presence of several local optima, in different regions of the parametric space. We search for $\max_{\delta\in\mathcal{D}}\;\alpha(A(K_{1},\delta))$.
@@
We ran the Sequential Quadratic Programming (SQP) algorithm from 100 random initial points. As termination criteria, we set $10^{-6}$ as optimality tolerance (KKT stationarity condition), and $10^{-8}$ as step tolerance (step size between two iterations). The resulting worst-case spectral abscissas $\alpha(A(\delta^{*}_{\text{SQP}}))=-a(\delta^{*}_{\text{SQP}})$ of each run are in Fig. 3(a), and suggest the presence of several local optima, which highlights the limitation of such a local method. These various local optima may correspond to significantly different parametric configurations. For example, the worst case 0.135 has the value -1.0 for the (normalized) value of one of the moments of inertia of the solar panel, -0.09 and 1.0 for the modal participation factors of the clamped-free modes 1 and 3 in rotation around the same axis; in contrast, the local optimum 0.007 has -0.30, -1.0 and -0.87 respectively for these three parameters. Excluding the runs that did not detect instability (each of which took less than 0.1s), the remaining runs took between 2.5 and 78s (average: 21s – regarding total computation time, we recall that each SQP run is inherently sequential, but that the 100 runs are executed in parallel).
@@
(b) MC and MC+SQP, repeated 100 times
@@
b) MC, MC + SQP
As explained in Section III-C2, and to compare the results with SQP alone, we evaluated 100 groups of 11000 samples, and the worst case in each group was selected to initialize an SQP run. Fig. 3(b) presents the worst case for each group (in blue) and the worst case after the subsequent SQP (in red). Overall, MC+SQP performs signicantly better than SQP alone or MC alone. Each group of 11000 samples, computed in parallel over 14 cores, took about 6.2s; the SQP runtimes were between 40s and 100s.
We mention that 78% of the drawn samples were not admissible ($c(\delta)>0$) and were therefore discarded before evaluation of $a(\delta)$. We also mention that the total number of runs $N_{\text{MC}}$ verifies $N_{\text{MC}}\approx\textrm{ln}(\gamma)/\textrm{ln}(1-\epsilon)$ with the values $\gamma=\epsilon=$1e-5, which guarantees as in Eq. (21) that the worst-case $a(\delta^{*}_{\text{MC}})$ found through all samples of Fig. 3(b) (without considering the subsequent SQP) covers 99.999% of the probability distribution of $a(\delta)$ with a confidence level of 99.999%. These observations suggest that the worst case around 0.136 is rare, even though unstable configurations are frequent (around 92% of the admissible configurations).
@@ V-D1 Implementation details
Compared to the choice of parameters of Section V-C, we reduced the step tolerance from 1e-8 to 1e-6 for the SQP algorithm. The objective is to speed up the detection of active configurations in most iterations, relying on Step 4 to perform a finer validation. Below, we present two different strategies for the worst-case searches.
Strategy 1 (single KKT point)
We run a single instance of MC+SQP, i.e. a single run of SQP chosen from the worst case among the $N$ MC samples. Although, as seen in Section V-C, this does not always converge to the global optimum, we prioritize computational speed by stopping at the first KKT point for several reasons (any missed problematic configuration may still be found in subsequent steps or iterations). The resulting (single) KKT point, if problematic, is added to the active configurations.
Strategy 2 (multiple KKT points)
In Strategy 2, we perform 4 instances of MC+SQP and obtain 4 KKT points $\delta^{*}_{i}$. For clarity, let us assume that they are all problematic configurations. The worst case among those 4, noted $\delta^{*}_{1}$, is added to the active configurations. Then, we compute the distance $|\delta_{2}^{*}-\delta_{1}^{*}|$; if the distance exceeds a specified threshold, $\delta^{*}_{2}$ is also added to the active configurations. The procedure is repeated for $\delta_{3}^{*}$ and $\delta_{4}^{*}$: each $\delta_{i}^{*}$ is added to the active scenarios only if its distance from all previously selected points exceeds the threshold. The threshold is chosen as half of the hypercube diagonal of the parameter space, i.e., $\sqrt{k}$, where $k$ is the number of uncertain parameters. This choice limits the number of active scenarios by enforcing a minimum spatial separation, thereby ensuring that only significantly distinct worst-case configurations are retained. Indeed, the computation time of the multi-models controller synthesis (Step 1) increases with the number of active configurations.
@@
In Step 4, we run a similar worst-case search as in Section V-C. We generated only 20 groups of MC samples (with one subsequent SQP run for each of the 20 groups), which is deemed sufficient for our application based on the results of Section V-C. This procedure is applied for both the stability and performance, and the Algorithm 2 terminates when no problematic configuration is found. In the implementation, we conducted the worst-case search for performance before that for stability to avoid unnecessary computations, since performance was observed to fail more frequently than stability.
@@ V-D2 Results
As in section V-C, the tests are ran on a 12th Gen Intel(R) Core(TM) i7-12700H 2.30 GHz, with 14 cores and 16GB of RAM, and computation times are provided for reference. We used the tolerances $\epsilon_{1}=5\%$ and $\epsilon_{2}=1\%$ (cf. Eq. (23)).
The Algorithm 2 was executed 5 times for each strategy. All runs converged to a robust controller. Table I presents a few averaged metrics for comparison. The runtime of the final validation (Step 4) is averaged over all 10 runs, because this step is conducted identically for each strategy. Fig. 4 presents the transfer functions of Problem (26) after optimization of the controller. Active configurations (in red) include the worst cases detected in the final iteration.
@@
Metrics Strategy 1 Strategy 2
Nb of iterations ( = nb. of executions of Step 1) (min/avg/max) 7/8.4/10 6/7.4/10
@@
Total runtime, excluding final validation (avg) 4034s 3708s
… including % represented by controller tuning (Step 1) 34% 33%
… including % represented by worst-case searches (Steps 2 and 3) 66% 67%
@@ VI Conclusion
This paper focused on the robust control and worst-case search in presence of nonlinear constraints restricting the hypercube. This specificity is missing from the current state of the art, and arises for example in mechanical systems.
@@
To perform the worst-case search, we combined Monte-Carlo sampling for global exploration and Sequential Quadratic Programming for local exploitation. Numerical experiments on a flexible spacecraft benchmark show that this method is able to reliably detect even rare worst-case configurations. Then, by iteratively constructing a set of active configurations, multi-objectives robust controller optimization was performed in reasonable runtime.
Finally, we emphasize that the proposed worst-case search, performing global exploration (albeit without deterministic guarantee of finding the global optimum) and being compatible with a large number of uncertain parameters and states (as a gradient-based method), has the potential to complement existing robust control techniques (systune, $\mu$-analysis) even in the unconstrained case.